### v1

# Data-Driven Discovery of Mobility Periodicity for Understanding Urban Transportation Systems

Abstract
Uncovering the temporal regularity of human mobility is crucial for discovering urban dynamics and has implications for various decision-making processes and urban system applications. This study formulates the periodicity quantification problem in complex and multidimensional human mobility data as a sparse identification of dominant positive auto-correlations in time series autoregression, allowing one to discover and quantify significant periodic patterns such as weekly periodicity from a data-driven and interpretable machine learning perspective. We apply our framework to real-world human mobility data, including metro passenger flow in Hangzhou, China and ridesharing trips in New York City (NYC) and Chicago, USA, revealing the interpretable weekly periodicity across different spatial locations over past several years. In particular, our analysis of ridesharing data from 2019 to 2024 demonstrates the disruptive impact of the COVID-19 pandemic on mobility regularity and the subsequent recovery trends, highlighting differences in the recovery pattern percentages and speeds between NYC and Chicago. We explore that both NYC and Chicago experienced a remarkable reduction of weekly periodicity in 2020, and the recovery of mobility regularity in NYC is faster than Chicago. The interpretability of sparse autoregression provides insights into the underlying temporal patterns of human mobility, offering a valuable tool for understanding urban systems. Our findings highlight the potential of interpretable machine learning to unlock crucial insights from real-world mobility data.
Keywords Human mobility  $\cdot$ Transportation systems  $\cdot$ Regularity  $\cdot$ Periodicity  $\cdot$ Time series  $\cdot$ Machine learning  $\cdot$ Optimization
1 Introduction
Human mobility in cities often reveals recurring patterns—the predictable ebbs and flows of people moving through urban areas—that reflect the underlying structure of urban life (Gonzalez et al., 2008; Song et al., 2010; Simini et al., 2012). From daily commutes to weekend trips, these temporal regularities in human mobility underpin everything from urban transportation management (Sheffi, 1985) and transit scheduling (Guihaire and Hao, 2008) to predicting the spread of diseases (Balcan et al., 2009; Belik et al., 2011; Castillo-Chavez et al., 2016; Li et al., 2022; Du et al., 2018) and optimizing resource allocation (Katoh et al., 2024). In essence, human mobility regularity is an important concept for identifying periodic trips in urban areas, which often follow daily and weekly cycles. A clear example is daily commuting, where passenger flows typically exhibit morning and afternoon peaks, forming a roughly “M"-shaped pattern in time series data (Chen and Sun, 2022) (see e.g., Fig. 1B). As cities evolve, mobility patterns have become increasingly complex and less predictable, characterized by complicated regulatory policies, diverse transportation modes and behavioral shifts driven by social economic and environmental factors (Midgley, 2009; Machado et al., 2018; Tirachini, 2020; Nouvellet et al., 2021; Li et al., 2022). Thus, understanding the rhythm of human mobility in urban areas is essential for grasping how our societies function and urban systems operate (Gonzalez et al., 2008; Song et al., 2010; Simini et al., 2012; Batty et al., 2012; Jiang et al., 2017; Alessandretti et al., 2020).
A paradigm for discovering long-term human mobility relies on collecting and constructing time series with a certain time resolution (e.g., hourly resolution) across different spatial locations over different years, which are often characterized as an algebraic tensor structure in Fig. 1A. In practice, tensors are multidimensional arrays which have been widely used to represent time series data of human mobility (Chen and Sun, 2022). To explore human mobility regularity underlying these time series, a few critical but unsolved questions arise as follows. First, there is no standard method to quantify and compare the strength of mobility regularity and periodicity with different noise levels and temporal structures. The two example time series of passenger flow in Figs. 1B-C show different periodic patterns, while the strength of periodicity can be intuitively revealed by the auto-correlation analysis in Fig. 1D. Although it is obvious to observe from Fig. 1D that the weekly periodicity of Fig. 1B is stronger than Fig. 1C, the question that remains is how to compute the comparable metric for time series periodicity by referring to the auto-correlation plots in Fig. 1D. Second, there is no established approach to leverage periodicity strength as a metric for tracking the spatial and temporal evolution of mobility systems. Previous research has adopted methods such as Fourier Transform (Li et al., 2010; Prabhala et al., 2014), entropy-based measures (Goulet-Langlois et al., 2017; Huang et al., 2019; do Couto Teixeira et al., 2021; Teixeira et al., 2021), reinforcement learning (Tao et al., 2021), and clustering of similar patterns (Manley et al., 2018; Zhong et al., 2015; Zhang and Song, 2022) to identify periodicity and regularity in human mobility. In dynamical systems, dynamic mode decomposition identifies the cyclical patterns through the eigenvalue decomposition of vector autoregressive coefficients (Brunton and Kutz, 2022).
However, these approaches typically do not provide an explicit and interpretable metric that directly quantifies periodicity, making them less suitable for downstream modeling or individualized interpretation. In addition, most studies that quantify periodicity or regularity in human mobility data focus on the analysis of individual trajectory data rather than aggregated demand or travel-mode usage data (Prabhala et al., 2014; Goulet-Langlois et al., 2017; Huang et al., 2019; Teixeira et al., 2021; do Couto Teixeira et al., 2021; Tao et al., 2021; Manley et al., 2018; Oliveira et al., 2016). While individual data can provide useful information on predicting mobility flows, interpretable information on aggregated demand data reveals longer term trends that can be more directly related to larger economic, social, and other changes that can help inform the development of policies in the future. Another fact is that mobility flows can be severely disrupted by extreme events and may evolve gradually over time due to changes in behavior, infrastructure, or policy. For example, the COVID-19 pandemic disrupted regular travel routines and led to a noticeable decline in mobility regularity (Li et al., 2022; Chen et al., 2025a). The challenge is how to systematically use periodicity strength to detect, quantify, and interpret such changes across both space and time.
Figure 1: Human mobility data and their time series periodicity. (A) Human mobility data can be structured as algebraic tensors where each entry corresponds to the value of a specific variable $\gamma$ at a particular spatial location $n$ and time step $t$. Here, $N$ represents the number of spatial locations, and $\delta$ refers to the number of variables or time phases. The variable dimension $\gamma$ is flexible and can encode different aspects of mobility, such as inflow and outflow at a particular spatial location. Moreover, the temporal lags such as data from different years can also be incorporated as distinct variables within this tensor framework. (B) Illustration of weekly periodicity in passenger flow time series. This panel displays an example of an hourly passenger flow time series that exhibits a clear weekly periodicity. This periodicity is visually highlighted by annotating the data points according to weekly cycles (e.g., blue and orange circles). The time series reveals distinct mobility patterns between weekdays and weekends, with weekdays showing more pronounced morning and afternoon peak hours. (C) Example time series of passenger flow for comparison. (D) Auto-correlation analysis through scatter plots. These panels present scatter plots that visualize the relationship between trip counts at a given time step $t$ and trip counts at previous time steps $t-k$, in which $k$ represents different time lags (e.g., $k=1,24,168$ correspond to hourly, daily, and weekly cycles, respectively). Points that align closely along the anti-diagonal in these scatter plots indicate a strong similarity between the trip counts at the two time steps, implying positive auto-correlation and potential periodicity. Notably, the scatter plot for a weekly cycle demonstrates a more remarkable similarity compared to the daily and hourly cycles, visually confirming the strong weekly periodicity in the example time series.
Substantial progress has been made in understanding human mobility regularity and patterns from analytical perspectives such as entropy and predictability (Oliveira et al., 2016; Barbosa et al., 2018; Delussu et al., 2023). However, the existing methods are not well-suited to quantifying periodicity across spatiotemporal domains and discovering dynamic patterns that related to human mobility regularity. Recently, interpretable machine learning methods (Murdoch et al., 2019; Rudin et al., 2022; Brunton and Kutz, 2022) offer new opportunities to improve appearance, robustness, and interpretability of data-driven exploration in mobility data. In terms of model development, sparse linear regression (Tibshirani, 1996; Jenatton et al., 2011) can be used to reformulate the problem of sparse identification of nonlinear dynamical systems (Brunton et al., 2016), which laid the foundations of several machine learning frameworks that designed for variable and feature selection in scientific discovery. In that venue, sparsity is one of the most important aspects to reinforce interpretability of machine learning models (Murdoch et al., 2019; Rudin et al., 2022). However, existing sparse methods face critical limitations when applied to large-scale human mobility data. First, classical sparse regression methods are primarily designed for variable selection from univariate and multivariate data and struggle to scale to multidimensional data with spatial and temporal dependencies. Second, mobility data are inherently noisy, irregular, and heterogeneous across spatiotemporal domains, which challenges standard sparse models that lack built-in mechanisms for structured robustness. Third, even when sparsity in enforced, the resulting models do not yield interpretable or comparable metrics for periodicity strength across different spatial locations, variables, and time phases.
To bridge these gaps, we propose a multidimensional sparse autoregression framework that models time series of mobility flows across space and time, explicitly captures dominant auto-correlations, and provides a unified and interpretable measure of periodicity strength. This interpretable machine learning framework builds upon the tensor data structure. Formally, human mobility data are spatially- and temporally-varying, and can be represented as a third-order tensor of non-negative integers $x_{n,\gamma,t}\in\{0,1,2,\cdots\}$ as shown in Fig. 1A, where $n\in[N]$ indexes spatial locations, $\gamma\in[\delta]$ indexes different variables or time phases, and $t\in[T_{\gamma}]$ indexes time steps. This algebraic tensor structure captures the multidimensional nature of mobility systems and aligns with a rich body of work on tensor-based machine learning for spatiotemporal pattern discovery (Kolda and Bader, 2009; Chen et al., 2024). Each individual sequence $x_{n,\gamma,t}$ is a time series, and thus can be seamlessly characterized by using time series models such as autoregression (Hamilton, 2020). As exemplified by Fig. 1B, real-world passenger flow time series demonstrate weekly periodicity by picking data points with a weekly cycle. When these data points are denoted by $x_{t},t\in[T]$, the auto-correlations in Fig. 1D can be formulated as $x_{t}\approx w_{k}x_{t-k}$, measuring self-similarity in mobility over any time lag $k$. Observing Fig. 1D, data points of the example time series are very close to the anti-diagonal curve at a weekly cycle (i.e., $k=168$), implying that weekly periodicity is positively auto-correlated. Moreover, the “closeness" of data points to the anti-diagonal at index $k=168$ in Fig. 1D justifies the strength of weekly periodicity of time series, demonstrating that the time series in Fig. 1C is less periodic at a weekly cycle than Fig. 1B.
2 Machine Learning Formula
Figure 2: Illustration of time series autoregression and its multidimensional setting on human mobility data. (A) The $d$th-order autoregression on the univariate time series can be formulated as an optimization problem for finding the optimal coefficients. Here, the order $d$ is an integer, determining the number of auto-correlations. (B) Multidimensional time series autoregression. Note that the symbol “$\times$" simply refers to the element-wise product between matrices. (C) The interpretability of time series autoregression can be reinforced by sparse and non-negative coefficients. The upper bound of the number of nonzero entries in the coefficient vector is set as an integer $\tau$, which is formally defined as the sparsity level. In mathematics, the cardinality of index set $\Omega$ (i.e., indices of nonzero coefficients) can be used to reformulate sparsity constraints. As a result, the decision variables in the constructed optimization problem include both coefficients and index set.
Interpretable machine learning provides a way to understand underlying patterns of complicated data and systems (Brunton et al., 2016; Murdoch et al., 2019). This study starts from the core idea of autoregression on time series, namely, the current value of the time series can be predicted as a linear combination of its past values (Hamilton, 2020). The objective is to find a set of coefficients that minimizes the sum of squared autoregressive errors (Fig. 2A) with predefined constraints such as sparsity and non-negativity on the coefficients as shown in Fig. 2C. Sparsity, enforced by the constraint $|\Omega|\leq\tau$, limits the number of nonzero coefficients to a small integer $\tau$. This implies that only a few past time lags significantly contribute to predicting the current value. For instance, Fig. 1D shows that the trip counts at time $t$ are highly consistent with the trip counts at time $t-168$, revealing predictable and periodic patterns.
In the context of human mobility periodicity, we expect dominant nonzero coefficients that uncover periodic patterns such as a 168-hour cycle for weekly periodicity. Non-negativity of the coefficients ensures that the identified auto-correlations are positive, aligning with an intuitive understanding of periodicity as similar patterns repeating over time (see e.g., Fig. 1D). Multidimensional time series autoregression follows the specification of auto-correlations in which the coefficients over $N$ spatial locations and $\delta$ variables can be optimized (Fig. 2B). The optimization explicitly aims to find a small set of relevant past time lags in the index set $\Omega$ over $N\times\delta$ coefficient vectors. Formally, on the human mobility data $\{x_{n,\gamma,t}\}_{n\in[N],\gamma\in[\delta],t\in[T_{\gamma}]}$, we formulate the multidimensional sparse autoregression as follows,
$\min_{\{{w}_{n,\gamma,k}\},\Omega}$ $\sum_{n=1}^{N}\sum_{\gamma=1}^{\delta}\sum_{t=d+1}^{T_{\gamma}}\Bigl{(}x_{n,\gamma,t}-\sum_{k\in\Omega}w_{n,\gamma,k}x_{n,\gamma,t-k}\Bigr{)}^{2}$ (1)
[…]
where $w_{n,\gamma,k},n\in[N],\gamma\in[\delta],k\in[d]$ are the non-negative coefficients of this $d$th-order autoregression formula, allowing one to quantify the positive auto-correlations such as daily and weekly periodicity as the dominant ones in time series. The last constraint is the normalization on the non-negative coefficients, letting the sum of the coefficients be one. Since the optimization problem is equivalent to the $\ell_{0}$-norm induced sparse autoregression, it can be addressed by mixed-integer programming solvers with exact solutions (Bertsimas et al., 2016, 2020, 2024; Tillmann et al., 2024). On such multidimensional tensor time series, we can discover the weekly periodicity of human mobility across different spatial locations and variables (or time phases).
On the hourly time series, if one intends to identify weekly periodicity, then the order of sparse autoregression should be set as $d\geq 168$ or greater. In what follows, we consider the order $d=168$ for the real-world mobility datasets. While the constraint $|\Omega|\leq\tau$ requires selecting a subset of $\tau$ coefficients out of $d$ possible ones, we extract the coefficients at index $k=168$, namely, auto-correlations $w_{n,\gamma,168},n\in[N],\gamma\in[\delta]$ to interpret the strengths of weekly periodicity across $N$ spatial locations over $\delta$ variables. The sparsity level $\tau$ is a critical hyperparameter for sparse autoregression. If this value is very small, then the index $k=168$ might not be in the support set $\Omega$, making it difficult to quantify weekly periodicity. In contrast, given a large sparsity level $\tau$, the values of auto-correlations within the support set $\Omega$ would have marginal differences.
3 Metro Passenger Flow Periodicity
To demonstrate the effectiveness of our framework in capturing periodicity from real-world mobility data, we begin with a case study of metro passenger flow data in Hangzhou, China. This dataset offers a clean and well-structured setting with strong known periodic patterns, making it an ideal testbed for validating the ability of our sparse autoregression model to recover meaningful temporal cycles—particularly weekly periodicity across different stations and flow directions.
Figure 3: Identification of weekly periodicity in inflow and outflow time series from the Hangzhou metro passenger flow dataset (from January 2 to 25, 2019) with the multidimensional sparse autoregression method (sparsity level $\tau=4$). (A) Daily average inflow trips of 81 metro stations. Note that the trip records of the station #54 are missing. (B) Weekly periodicity of hourly inflow time series achieved by the sparse autoregression with sparsity level $\tau=4$. 64 stations show high periodicity ($\geq$ 0.8). The average periodicity value is 0.844. (C) Weekly periodicity of hourly outflow time series. 53 stations show high periodicity ($\geq$ 0.8). The average periodicity value is 0.829. Since inflow and outflow are two variables in the tensor data, their coefficients are optimized simultaneously with the same index set. (D) Scatter plots of trip counts at time $t$ versus time $t-168$ of 6 selected stations, illustrating varying strengths of weekly periodicity. The time series with a high weekly periodicity value also presents a strong similarity of data points between times $t$ and $t-168$. (E) Hourly inflow time series for stations #14 and #15, shown across three consecutive weeks: January 2–8, January 9–15, and January 16–22. Strong overlap among these weekly segments indicates high periodicity at a weekly cycle.
[…]
Table 1: Strength of weekly periodicity of the Hangzhou metro passenger flow data at the selected end stations of metro lines.
Station # Inflow Outflow Station # Inflow Outflow
23 0.77 0.65 35 0.91 0.76
24 0.46 0.48 36 0.88 0.74
25 0.24 0.54 37 0.85 0.74
26 0.78 0.62 38 0.76 0.75
27 0.93 0.78 39 0.73 0.74
28 0.87 0.69 40 0.83 0.88
29 0.77 0.66 62 0.95 0.82
30 0.81 0.71 63 0.89 0.77
31 0.90 0.70 64 0.89 0.81
32 0.91 0.74 65 0.86 0.80
33 0.90 0.80 66 0.87 0.73
34 0.90 0.76 67 0.93 0.79
Fig. 3D shows the weekly periodicity values for six selected stations, measured as the strength of auto-correlations at a weekly lag (i.e., $k=168$) in the inflow time series. The source emphasizes that time series with high weekly periodicity values also show a strong similarity between data points at times $t$ and $t-168$, indicating positive auto-correlations. The weekly periodicity is visually confirmed by scatter plots and time series overlaps. For instance, station #14 shows a high weekly periodicity of 0.93, and its scatter plot shows a strong alignment along the anti-diagonal (Fig. 3D). In particular, the comparison between stations #14 and #15 in Fig. 3E answers why the strength of time series periodicity is a fair metric to reveal repeating patterns of human mobility. The periodicity value of station #15 is 0.71, which is smaller than station #14. As shown in Fig. 3E, the trip count time series at station #15 does not align perfectly across weeks, indicating weaker weekly periodicity compared to the more consistent pattern observed at station #14. To summarize, on the passenger flow dataset, the comparable metric of weekly periodicity across different stations and inflow/outflow variables to discover periodic mobility patterns in a city-wide scale.
4 Regularity of Ridesharing Trips
Figure 4: Identification of weekly periodicity in ridesharing trip data across different urban areas with the multidimensional sparse autoregression method (sparsity level $\tau=6$). (A) Weekly periodicity of NYC ridesharing pickup trip data across 265 areas from 2019 to 2024. The average periodicity values are also marked. (B) Weekly periodicity of Chicago ridesharing pickup trip data across 77 areas from 2019 to 2024. (C) The coefficient difference between 2024 and 2019, referring to the coefficients of 2024 minus 2019. (D) Annual ridesharing trip counts of NYC and Chicago from 2019 to 2024. In contrast to NYC, the ridesharing trips in Chicago show a poor recovery in the post-pandemic years. Note that the ridesharing trip data of January 2019 is not available from the data portal. (E) Weekly average trip time series of NYC ridesharing, starting from Monday to Sunday. (F) Weekly average trip time series of Chicago ridesharing, starting from Monday to Sunday. While NYC shows the recovery percentage as 68% in 2021, Chicago keeps 44%, roughly as the same level of 2020.
Multidimensional sparse autoregression quantifies human mobility periodicity through coefficients. The resulting non-negative coefficients directly quantify the strength of the positive auto-correlation at certain time lags. A high coefficient at a time lag corresponding to a weekly cycle (e.g., 168 hours for hourly time series data) indicates a strong weekly periodicity.
As we know, the COVID-19 pandemic had a significant disruptive impact on urban mobility regularity and periodicity. To examine how such disruptions affected periodicity, we analyze large-scale ridesharing trip data from 2019 to 2024 in NYC [^2] The NYC ridesharing trip record dataset is available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page. and Chicago [^3] The Chicago ridesharing trip record dataset is available at https://data.cityofchicago.org/.. This analysis aims to demonstrate how the strength of weekly periodicity varied across different urban areas and years, reflecting the impact of the pandemic. In the experiments, we set the sparsity level and the order of sparse autoregression as $\tau=6$ and $d=168$, respectively. On both NYC and Chicago ridesharing datasets, the support set is optimized as $\Omega=\{1,23,24,143,167,168\}$, covering the indices of local auto-correlations, daily cycle, and weekly cycle.
In NYC, the weekly periodicity of pickup trips in 2019 was similar to the patterns observed in 2022–2024, indicating a return to pre-pandemic regularity after disruptions (Fig. 4A). The year 2020 marked a significant shift from this regularity in NYC, with the pandemic severely disrupting trip patterns and reducing overall periodicity strength. By 2021, the ridesharing trips in NYC began recovering, transitioning from pandemic-induced irregularity towards the post-pandemic norm. During this recovery phase, suburban areas displayed increasingly periodic trip patterns in NYC. The periodicity patterns of ridesharing trips in 2022 became consistent with the patterns in 2019, while both periodicity values are almost same. Across these 6 years, ridesharing trips in downtown areas such as Manhattan were less periodic than suburban areas. In 2024, the overall weekly periodicity of NYC has been reinforced slightly. The ridesharing trips in suburban areas tend to be more periodic than 2019, while the downtown areas were less periodic (Fig. 4C).
In Chicago, a similar pattern emerged where the overall periodicity of ridesharing trips in 2019 was stronger than in 2020, and downtown areas in 2020 were less periodic than in 2019 (Fig. 4B). However, in contrast to NYC, the periodicity pattern transition in Chicago was slower, with no remarkable pattern change between 2020 and 2021. The periodicity patterns of ridesharing trips in Chicago during 2022–2024 also became consistent with patterns in 2019, suggesting a return to pre-pandemic regularity. In 2024, although the ridesharing trips demonstrated the same level of weekly periodicity as 2019, downtown areas became less periodic (Fig. 4C). Comparing the aggregated trip counts in Fig. 4D, NYC showed a quick return of ridesharing trip counts and reached the same level as 2019, while Chicago showed a poorer recovery in the post-pandemic years. The weekly average trip time series also indicated that while NYC showed a recovery percentage of 68% in 2021 (Fig. 4E), Chicago remained at 44%, similar to its 2020 level (Fig. 4F). Although Chicago showed a recovery percentage of 76% ridesharing trips in 2024 (Fig. 4F), the weekly periodicity already reached the same level as 2019.
Our analysis so far suggests that the COVID-19 pandemic led to a significant decrease in the periodicity of urban mobility in both NYC and Chicago in 2020. Both cities showed signs of recovery in subsequent years, with periodicity patterns returning to pre-pandemic levels by 2022-2024. However, the pace of recovery in terms of both periodicity and overall trip counts appeared to be faster in NYC compared to Chicago. Furthermore, our method enables policymakers to measure disruptions such as pandemics and design targeted recovery strategies.
5 Discussion
Understanding the periodicity of human mobility is critical for revealing the underlying structure of individual and collective movement patterns. Periodic patterns in human mobility—ranging from daily commutes to weekly rituals—are manifestations of deeply rooted behavioral rhythms. These patterns are essential components for making informed decisions, developing effective urban policies, and building more resilient and responsive urban systems. In literature, such rhythms have been extensively studied in chronobiology, behavioral sciences, and temporal network theory (Roenneberg et al., 2004; Dijksterhuis et al., 2006; Holme and Saramäki, 2012), where periodicity reveals regularities in decision-making, attention cycles, and social synchronization. Drawing on these insights, we propose a novel method for quantifying and leveraging periodicity in large-scale mobility data. This enables a deeper understanding of urban rhythms and enhances predictive modeling for transportation planning, epidemic spread (Balcan et al., 2009; Belik et al., 2011; Castillo-Chavez et al., 2016), and urban resilience assessments (Xu et al., 2025).
Our results advance people’s understanding of real-world urban systems and show the capability of multidimensional sparse autoregression for quantifying mobility regularity from time series data. Our findings concerning the weekly periodicity—a comparable metric across different time series, spatial areas, and time periods—in NYC and Chicago reveal the disruptive impact of the COVID-19 pandemic on mobility regularity and subsequent recovery trends. The interpretability offered by the multidimensional sparse autoregression is a key advantage, as sparse and non-negative auto-correlations provide actionable insights into the specific temporal lags driving the observed periodicity (Chen et al., 2025b). The method we present allows substantial analysis of time series data that collected from real-world systems and demonstrated periodicity or seasonality. Despite identifying weekly periodicity, the revealed dominant auto-correlations in the support set enable us to introduce proper differencing operations—a simple yet effective solution to address the non-stationarity issue (Chen et al., 2025c).
As the field of human mobility research moves from analytical studies to data-driven experiments designed to directly learn underlying patterns, there is a growing need to formulate the periodicity quantification problem as a sparse identification of dominant positive coefficients in multidimensional time series autoregression. We have developed an interpretable machine learning method for identifying spatial patterns and temporal dynamics that relate to weekly periodicity in urban areas. While this study focuses on metro passenger flow and ridesharing trips in specific cities, a potential limitation is the focus on these modes. Future work could integrate data from other transportation modes, such as public transit, to capture more comprehensive multi-modal behaviors. In terms of potential applications, this interpretable framework provides a robust foundation for discovering and quantifying temporal patterns related to periodicity and seasonality in complicated data and systems beyond human mobility, including climate systems and healthcare time series.
Appendix A Data, Material, and Code Availability
Data and code of this study will be available at https://github.com/xinychen/integers. For an intuitive understanding of time series periodicity, we also provide an interactive visualization tool on the ridesharing trip time series.
Appendix B Methodology
While the optimization problem we have posed has the potential to learn a significant amount of information about periodicity in time series data, estimating the coefficient vectors $\{{w}_{n,\gamma,k}\}_{n\in[N],\gamma\in[\delta],k\in[d]}$ of $N$ spatial areas and $\delta$ phases when $N$ or $\delta$ become even moderately large becomes extremely computationally expensive. Thus, instead of estimating individual coefficient vectors $\{{w}_{n,\gamma,k}\}_{n\in[N],\gamma\in[\delta],k\in[d]}$, we first estimate global auto-correlations with sparsity constraints. The optimization problem for estimating a global coefficient vector $\boldsymbol{w}=\{w_{k}\}_{k\in[d]}\in\mathbb{R}^{d}$ from the human mobility data $\{x_{n,\gamma,t}\}_{n\in[N],\gamma\in[\delta],t\in[T_{\gamma}]}$ is given by
$\min_{\{w_{k}\},\Omega}$ $\sum_{n=1}^{N}\sum_{\gamma=1}^{\delta}\sum_{t=d+1}^{T_{\gamma}}\Bigl{(}x_{n,\gamma,t}-\sum_{k\in\Omega}w_{k}x_{n,\gamma,t-k}\Bigr{)}^{2}$
[…]
which can be solved by mixed-integer programming. The support set is therefore optimized $\Omega$ simultaneously.
Given the global support set $\Omega$, learning the individual coefficients $\{w_{n,\gamma,k}\}_{n\in[N],\gamma\in[\delta],k\in[d]}$ can be simplified as follows,
$\min_{\{w_{n,\gamma,k}\}}$ $\sum_{t=d+1}^{T_{\gamma}}\Bigl{(}x_{n,\gamma,t}-\sum_{k\in\Omega}w_{n,\gamma,k}x_{n,\gamma,t-k}\Bigr{)}^{2}$
[…]
for any $n\in[N]$ and $\gamma\in[\delta]$. We solve this optimization problem by using quadratic programming.
Appendix C Metro Passenger Flow Periodicity
C.1 Weekly periodicity on the 30-minute passenger flow time series data
Fig. LABEL:Hangzhou_k336 shows the weekly periodicity across 80 stations by using the proposed model. While this study discusses passenger flow time series with both hourly and 30-minute time resolutions, it is possible to see the periodicity pattern shift of passenger flow from hourly data to 30-minute data. On these time series, the sparsity level of the proposed model is set as $\tau=4$, and the support set is optimized as $\Omega=\{1,48,335,336\}$ (i.e., covering both daily cycle $k=48$ and weekly cycle $k=336$). Thus, the weekly periodicity is revealed by the auto-correlation at the index $k=336$. Comparing to the hourly time series, the weekly periodicity in Fig. LABEL:Hangzhou_k336 is relatively small because 30-minute time series demonstrate stronger local auto-correlations than hourly time series. Of the results, the inflow is more periodic than the outflow in the end stations, while the inflow is less periodic than the outflow in downtown areas. Thus, the time resolution impacts the time series periodicity quantification. Since the periodicity is a relative value, the comparison among time series with the same time resolution is still meaningful.
Appendix D Regularity of Ridesharing Trips
D.1 Weekly periodicity of ridesharing pickup trips with different sparsity levels
In this study, we consider the sparsity level on the NYC ridesharing data as $\tau=6$, covering weekly cycle in the support set. For comparisons, we set the sparsity level of the proposed model as $\tau=4$ in which the support set is optimized as $\Omega=\{1,23,167,168\}$. As shown in Fig. 6, although the weekly periodicity at the sparsity level $\tau=4$ is a little different from $\tau=6$, we can also observe the reduction of weekly periodicity in 2020 due to the COVID-19 pandemic, following by the recovery of human mobility regularity to the normal status. Thus, the interpretability of weekly periodicity is robust to the setting of different sparsity levels.
Figure 6: Weekly periodicity of ridesharing pickup trips quantified by the proposed model with sparsity level $\tau=4$.
D.2 Weekly periodicity of ridesharing dropoff trips with different sparsity levels
In terms of dropoff trips, we set the sparsity levels of the proposed model as $\tau=4,6$ as follows.
  • The support set with sparsity level $\tau=4$ is $\Omega=\{1,23,167,168\}$.
  • The support set with sparsity level $\tau=6$ is $\Omega=\{1,23,24,143,167,168\}$.
As shown in Fig. LABEL:NYC_dropoff_periodicity_tau, the dropoff trips of NYC ridesharing data demonstrated similar temporal patterns of weekly periodicity as the pickup trips. Trips with destination to downtown areas are less periodic than suburban areas. In 2020, a remarkable reduction of weekly periodicity can be clearly observed. In 2022, the weekly periodicity recovered to the same level as 2019. These findings intuitively verify the impact of the COVID-19 pandemic on mobility regularity in urban areas.